. In the zwitterionic (charge separated, + and – formal charges present) resonance form of the amino acid glycine (side chain R = H), the distance from the positively charged amino nitrogen to the negatively charged oxygen atoms on the carboxylate, on the other end of the molecule, is about 3 Å. (a) Sketch a 3D structure, keeping in mind both of the relevant resonance structures, and determine the hybridization scheme used in the σ-bonding framework. (b) Calculate the dipole moment, = . Express your answer in Debye (D). Recall that 1 D = 3.34 × 10./0 C m. The net charge at the two ends of the molecule, and , should be converted to coulombs (C) using the elementary charge = 1.602 × 10.67 C, the charge on an electron or a proton. Use as the distance the distance between the atoms that carry the formal charges and .
Answer #1

Part a

The resonance structures of the zwitter ion is as follows.

3 3

Now for the hybridization, the negatively charged oxygen, the positively charged nitrogen and one carbon are sp3 hybridized, while the double bonded oxygen and the other carbon are sp2 hybridized.

sp3 sp2 sp2 H3N sp3 sp3

Part b

The dipole moment can be calculated as,

\mu =q*r

where q is the seperated charge and r is the distance. Susbstituting the values and converting them to the appropriate units, we have,

\mu =\frac{1.602*10^{19}C*3*10^{-10}m*1D}{3.34*10^{-30}Cm}=14.39D